OpenStudy (anonymous):

anyone good with integration by parts? Integrate ((x)e^(2x))/(1+2x)^2

OpenStudy (anonymous):

\[\int\limits_{}^{}\frac{ xe ^{2x} }{ (1+2x)^2 }\]

OpenStudy (anonymous):

ok ive gotten to where you are

OpenStudy (anonymous):

i mean understanding what you did

ganeshie8 (ganeshie8):

small typo, let me fix it quick

ganeshie8 (ganeshie8):

Let : \(u=xe^{2x} \implies du = (1\cdot e^{2x}+x\cdot 2e^{2x})dx = e^{2x}(1+2x)dx\) and \(dv = \dfrac{1}{(1+2x)^2} dx\implies v = -\dfrac{1}{2(1+2x)}\)

ganeshie8 (ganeshie8):

\[\begin{align}\int\dfrac{ xe ^{2x} }{ (1+2x)^2 }~~ &=~~ xe^{2x}\cdot \dfrac{-1}{2(1+2x)} - \int e^{2x}(1+2x)\cdot \dfrac{-1}{2(1+2x)} dx\\~\\ &=~~ xe^{2x}\cdot \dfrac{-1}{2(1+2x)} + \int \dfrac{e^{2x}}{2}dx \\~\\ \end{align}\]

ganeshie8 (ganeshie8):

rest should be easy is there any other function thats as easy as exponential function to integrate ?

OpenStudy (anonymous):

so i should let u = e^2x and du = e^2

ganeshie8 (ganeshie8):

are you saying you dont know how to integrate \(e^{2x}\) ?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

u =2x, du =2, dv = e , v =e?

ganeshie8 (ganeshie8):

ohkie, its easy, whats the derivative of \(e^{2x}\) ?

OpenStudy (anonymous):

e^(2x)?

ganeshie8 (ganeshie8):

heard of chain rule before ?

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

e^(2x)2

ganeshie8 (ganeshie8):

right, so an antiderivative of e^(2x)2 is e^(2x)

ganeshie8 (ganeshie8):

another name for "antiderivative" is "indefinite integral"

ganeshie8 (ganeshie8):

lets do few more examples so that you see how to to guess these

ganeshie8 (ganeshie8):

whats the derivative of \(\sin x\) ?

OpenStudy (anonymous):

cos x

ganeshie8 (ganeshie8):

so an antiderivative of \(\cos x\) is \(\sin x\) : \(\int \cos x\,dx = \sin x+C\)

OpenStudy (anonymous):

-sinx +c?

ganeshie8 (ganeshie8):

its +sinx because the derivative of \(\sin x\) is \(\cos x\), the antiderivative of \(\cos x\) is \(\sin x+C\)

OpenStudy (anonymous):

ok got it

ganeshie8 (ganeshie8):

lets do couple more find an antiderivative of \(\large x^2\)

OpenStudy (anonymous):

x^3/3

ganeshie8 (ganeshie8):

correct. but why ?

OpenStudy (anonymous):

the formula is x^n+1/n+1 ?

ganeshie8 (ganeshie8):

nope, don't use the formula use the fact that differentiating x^3/3 gives you x^2, so x^3/3 is an antiderivative of x^2

OpenStudy (anonymous):

ok

ganeshie8 (ganeshie8):

find an antiderivative of \(e^{2x}\)

OpenStudy (anonymous):

that makes sense ill just from now on relate that

ganeshie8 (ganeshie8):

in other words, what function do u need to differentiate in order to produce \(e^{2x}\)

OpenStudy (anonymous):

I was considering how: \[ \left(\frac{u}{v}\right)' = \frac{u'v-uv'}{v^2} \]I put \(v=1+2x\), and then got the equation: \[ (1+2x)u' - (2)u = xe^{2x} \]I divide both sides by \(x\) and got: \[ \frac{u'-2u}{x} +2u' = e^{2x} \]I came to the impression that if I can get \(u'=2u\), I could make things simpler. So I used \(u=Ce^{2x}\).\[ \frac{2Ce^{2x}-4Ce^{2x}}{x} +4Ce^{2x} = e^{2x} \implies 4Ce^{2x} =e^{2x} \implies C=\frac 14 \]That got me to my solution of\[ \frac uv = \frac{\frac{1}{4}e^{2x}}{1+2x} = \frac{e^{2x}}{4+8x} \]This isn't a reliable method, but I think it's one of the few cases where you can reverse engineer the quotient rule.

OpenStudy (anonymous):

i cant do it that way though thats cool

OpenStudy (anonymous):

(e^(2x))/2

OpenStudy (anonymous):

also wio that is the correct answer

OpenStudy (anonymous):

v=e^2/2

ganeshie8 (ganeshie8):

Nice! as you can see integration is indeed a guessing game and you can only get good at it by practice

ganeshie8 (ganeshie8):

are you saying (e^(2x))/2 is an antiderivative of e^(2x) ?

OpenStudy (anonymous):

yes

ganeshie8 (ganeshie8):

why

OpenStudy (anonymous):

cause the derivative of that is e^2x

ganeshie8 (ganeshie8):

Perfect! i wanted to hear that from you :)

ganeshie8 (ganeshie8):

so you can work the rest of the problem

ganeshie8 (ganeshie8):